### v1

# Lifelong Personalized Low-Rank Adaptation of Large Language Models for Recommendation

Abstract.
We primarily focus on the field of large language models (LLMs) for recommendation, which has been actively explored recently and poses a significant challenge in effectively enhancing recommender systems with logical reasoning abilities and open-world knowledge. Current mainstream efforts mainly center around injecting personalized information from recommendation models into LLMs by customizing input templates or aligning representations between semantic and recommendation spaces at the prediction layer. However, they face three significant limitations: (1) LoRA is mostly used as a core component in existing works, but personalization is not well established in LoRA parameters as the LoRA matrix shared by every user may not cater to different users’ characteristics, leading to suboptimal performance. (2) Although lifelong personalized behavior sequences are ideal for personalization, their use raises effectiveness and efficiency issues since LLMs require escalating training and inference time to extend text lengths. (3) Existing approaches aren’t scalable for large datasets due to training efficiency constraints. Thus, LLMs only see a small fraction of the datasets (e.g., less than 10%) instead of the whole datasets, limiting their exposure to the full training space. To address these problems, we propose RecLoRA. This model incorporates a Personalized LoRA module that maintains independent LoRAs for different users and a Long-Short Modality Retriever that retrieves different history lengths for different modalities, significantly improving performance while adding minimal time cost. Furthermore, we design a Few2Many Learning Strategy, using a conventional recommendation model as a lens to magnify small training spaces to full spaces. Extensive experiments on public datasets demonstrate the efficacy of our RecLoRA compared to existing baseline models.
Large Language Models, Recommender Systems, User Modeling
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data miningccs: Information systems Recommender systems
1. Introduction
Figure 1. The illustration of lifelong behavior sequence and personalized low-rank adaption (LoRA) of large language models for recommendation.
Recommender systems play essential roles in various online services to mitigate the information overload problem and meet users’ information needs (Guo et al., 2017a; Lin et al., 2023b; Liu et al., 2024; Qin et al., 2022, 2023). Besides, large language models (LLMs) have experienced significant growth in the field of natural language processing (NLP), demonstrating a remarkable ability to comprehend human language and generate text that closely resembles human writings for a broad range of tasks (Zhao et al., 2023; Brown et al., 2020; Touvron et al., 2023). Recent studies have started to investigate the potential of LLMs for recommender systems under a variety of recommendation tasks, such as listwise ranking and pointwise scoring (Lin et al., 2023a, 2024b). These studies usually inject personalized knowledge from recommendation domains into LLMs. (Lin et al., 2024b; Zhang et al., 2023b; Zheng et al., 2023). This personalized knowledge plays a pivotal role in tailoring experiences to individual preferences and enhancing user interactions and engagements (Chen et al., 2023). So our crucial points are what useful personalized knowledge we can provide and how to inject it into LLM to help with predictions.
As the proverb ”Actions speak louder than words” suggests, a user’s behavior sequence is the most direct expression of their preferences. It typically provides the most detailed personalized information about users in recommendation scenarios. As illustrated in Figure 1, if a user’s behavior can be collected and modeled from birth to the present, their interests can be accurately predicted. Several mainstream studies have designed memory mechanisms to model lifelong behavior sequences (Ren et al., 2019; Pi et al., 2019). Other approaches employ retrieval methods to extract relevant information from entire lifelong sequences (Qin et al., 2020; Pi et al., 2020a; Chang et al., 2023). These methods prove the effectiveness of user behavior sequences. However, they fail to leverage general knowledge of LLMs to help users’ predictions.
Regarding the second point, there are already existing studies on how to incorporate personalized knowledge into large language models (LLMs). Some works inject personalized information from conventional recommendation models (CRMs) into LLMs by customizing input templates at the input layer (Zhang et al., 2023b; Li et al., 2023b; Lin et al., 2024a; Bao et al., 2023). Other studies align representations between semantic and recommendation spaces at the prediction layer (Li et al., 2023a; Wang et al., 2023). Despite achieving promising results compared to conventional recommendation models like SASRec (Kang and McAuley, 2018a) and DIN (Zhou et al., 2018), existing works still face three significant limitations.
Firstly, personalization is not well established when adapting LLMs to recommender systems. For the sake of memory efficiency, mainstream approaches typically adopt low-rank adaptation (LoRA) techniques (Hu et al., 2021; Dettmers et al., 2023) and perform parameter-efficient finetuning (PEFT) to inject personalized knowledge from recommendation domains into LLMs (Lin et al., 2024b; Zhang et al., 2023b; Zheng et al., 2023). Ideally, each user should have a unique LoRA matrix to achieve personalized recommendations. However, current methods only implement prompt personalization or representation personalization, while the LoRA module, a core component of fine-tuning LLMs, remains static for all users and target items. This lack of parameter personalization results in inferior capabilities to effectively manage user dynamics and preference shifts in modern recommender systems.
Secondly, utilizing lifelong personalized behavior sequences in the prompt or representation layer raises issues of effectiveness and efficiency. Since LLMs are sensitive to input length and have a limit on input size, they tend to perform poorly on long behavior sequences, resulting in a ceiling on effectiveness.(Lin et al., 2024b) Additionally, as the length of the input behavior sequence and the volume of training data increase, the time required for training and inference escalates rapidly, leading to efficiency problems.
Thirdly, existing approaches are not scalable for large-scale industrial datasets due to training inefficiency. Even with LoRA techniques, fine-tuning LLMs on recommendation data consumes immense computational resources and time, making it extremely inefficient for large-scale datasets, typically involving millions or even billions of records. Although some studies (Lin et al., 2024b) suggest fine-tuning LLMs on a small fraction (e.g., less than 10%) of the training data to balance recommendation performance and training efficiency, this approach limits the LLMs’ exposure to the full training space, resulting in suboptimal performance. Therefore, enabling LLMs to fully perceive the training space while ensuring training efficiency remains an open research question.
To address the aforementioned issues, we propose a novel Personalized Low-rank Adaptation Framework of Large Language Models for Recommendation (i.e., RecLoRA). Specifically, we maintain a set of parallel, independent LoRA weights instead of a single static LoRA module as suggested in previous works (Lin et al., 2024b; Zhang et al., 2023b). As shown in Figure 1, this personalized LoRA module functions to incorporate personalized user knowledge into large language models, serving as a strategy for parameter personalization. We assign a sequential CRM and adapt the sequence representation to personalized meta-LoRA weights. More precisely, we use a soft routing method to aggregate meta-LoRA weights guided by the CRM. Instead of extending user behavior sequences in the text input, RecLoRA assigns longer sequences in the sequential CRM, which significantly improves effectiveness while adding minimal time cost. Moreover, the CRM is pre-trained on the full training space, so transforming the CRM to meta-LoRA weights acts like a lens, expanding the small fraction of training data exposed to the LLM to encompass the entire training space without increasing LLM training time.
The main contributions are summarized as follows:
  • To the best of our knowledge, we are the first to consider LoRA parameters personalization in the recommendation task. We design a personalized LoRA module that includes a CRM on full training space and a soft routing method to aggregate meta-LoRA weights with CRM instructions.
  • We propose a novel framework, RecLoRA, which realizes a better trade-off between effectiveness and efficiency. RecLoRA has a Few2Many Magnifier where only few-shot training data is needed to efficiently tune LLM and a Long-short Modality Retriever module, which extends the sequence length on the ID input side to improve performance while almost not increasing the time cost.
  • Our personalized LoRA module is an easy-to-plugin module, which can be used in any LLM4rec model with LoRA modules.
We have conducted offline experiments on three publicly available datasets and achieved promising results, demonstrating the effectiveness of our proposed model.
2. Preliminary
2.1. Problem Formulation
The core task of recommender systems is to estimate users’ preference toward target items given a certain context. We denote the historical dataset as $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $N$ is the number of data instances. The input $x_{i}$ contains features of user profile $u_{i}$(e.g., age, location, and historical behaviors), item attributes $v_{i}$ (e.g., category and brand), context information $c_{i}$ (e.g., time and season) and user history sequence $\mathcal{H}_{u_{i}}$(e.g., behaviors like click and ratings).
(1) $x_{i}$ $=(u_{i},v_{i},c_{i},\mathcal{H}_{u_{i}})$
where $\mathcal{H}_{u_{i}}=\{h_{1},h_{2},\dots,h_{N_{u_{i}}}\}$ represents $N_{u_{i}}$ sequential behaviors of user $u_{i}$.
The label $y_{i}\in\{1,0\}$ indicating the interaction signal of the user towards the item (i.e., prefer or not).
(2) $y_{i}=\begin{cases}1,~{}u_{i}~{}\text{clicks or likes}~{}v_{i};\\ 0,~{}\text{otherwise.}\end{cases}$
According to different input data transformations, there generally exist two different recommendation paradigms: (1) The input of ID modality $x_{i}^{ID}$ obtained by one-hot encoding for conventional recommendation models (CRMs), and (2) The input of textual modality $x_{i}^{text}$ generated by hard prompt template for LLMs as recommenders.
2.2. Conventional Recommendation Models
For ID input $x_{i}^{ID}$, various conventional recommendation models (CRMs) are designed to capture the collaborative patterns for precise user preference estimation from various aspects (e.g., feature interactions (Wang et al., 2021a; Guo et al., 2017a), user behavior modeling (Pi et al., 2020a; Qin et al., 2020)). We give the general formulation of CRMs as follows:
(3) $h_{i}^{ID}$ $=\operatorname{CRM}(x_{i}^{ID}),$
$\hat{y}_{i}^{ID}$ $=\sigma(\operatorname{MLP}(h_{i}^{ID}))\in(0,1),$
where $\sigma(\cdot)$ is the sigmoid function.
2.3. Large Language Models as Recommenders
Typically, large language models (LLMs) refer to Transformer-based language models with at least billion-level parameters that are trained on massive text datasets, demonstrating remarkable capacity in various natural language tasks. When adapting directly LLMs as the recommenders, we need to convert the raw data $(x_{i},y_{i})$ into textual input-output pairs $(x_{i}^{text},y_{i}^{text})$ with hard prompt template. We illustrate one template example in Figure 2.
Figure 2. The illustration of textual input-output pair.
As shown in Figure 2, the textual input $x_{i}^{text}$ consists of the user’s profile and historical behaviors, followed by a binary question about the personalized preference towards the target item. Following previous works (Lin et al., 2024b; Bao et al., 2023), to obtain the floating-point preference estimation $\hat{y}_{i}^{text}\in[0,1]$ instead of discrete word tokens, we apply bidimensional softmax over the corresponding scores of the binary key answer words (i.e., Yes & No) from LLMs to accomplish the preference estimation during evaluation:
(4) $\hat{y}_{i}^{text}=\frac{\exp(s_{i,\text{Yes}})}{\exp(s_{i,\text{Yes}})+\exp(s_{i,\text{No}})}\in(0,1).$
2.4. Low-Rank Adaptation of LLMs
Low-rank adaptation (LoRA) (Hu et al., 2021) serves as a popular parameter-efficient finetuning (PEFT) method to reduce the resource consumption of finetuning LLMs that possess massive parameters. The basic idea of LoRA is to maintain two trainable lightweight matrices $A$ and $B$ that are attached to a frozen pretrained weight matrix $W$. Hence, the linear transformation $Y=XW$ is reformulated as:
(5) $Y^{\prime}=XW+XAB^{T}=X(W+AB^{T}),$
[…]
(6) $A\sim\mathcal{N}(0,\sigma^{2}),\;B=0,$
so as to ensure the initial output $Y^{\prime}=XW+XAB^{T}$ equal to $Y=XW$. During finetuning, $W$ is fixed while $A$ and $B$ are updated by SGD-based optimization methods like Adam (Kingma and Ba, 2014).
When applying LoRA to LLM finetuning, we can selectively attach LoRA weights to target matrices of certain types inside the LLM. e.g., the matrices of query, key and value in the self-attention modules or the linear matrices in feed-forward networks (FFNs).
3. Methodology
In this section, we introduce our proposed RecLoRA (Personalized Low-rank Adaptation for Recommendation) framework and its training process in details.
Figure 3. The model structure and learning strategy of our proposed RecLoRA framework.
3.1. Overview of RecLoRA
We propose the RecLoRA framework to address three major limitations: (1) To construct personalized low-rank adaptation for LLMs in recommendation, we employ a parallel meta-LoRA mechanism incorporating personalized knowledge. (2) To alleviate the effectiveness and efficiency issues associated with behavior sequence extension, we design the Long-Short Modality Retriever. (3) To expand the receptive field of LLMs to the entire training space, we develop the Few2Many Magnifier Strategy. In this section, we provide a comprehensive explanation of our RecLoRA framework, including its network architecture and training process.
Figure 3 illustrates the overall framework of the RecLoRA model. As is shown in the figure, RecLoRA takes two input components: samples in ID modality $x^{ID}$ and samples in text modality $x^{text}$. Each modality $x$ has user profile, candidate item, context features and user history behaviors as formulated in Eq. (1). So we start by transforming the recommendation dataset into these two modalities and then feed $x^{ID}$ and $x^{text}$ into CRM and LLM models respectively. During transformation, a Long-Short Modality Retriever retrieves long history for $x^{ID}$ and short history for $x^{test}$.
(7) $x^{ID}$ $=LSMR(u^{ID},v^{ID})$
$x^{text}$ $=LSMR(u^{text},v^{text})$
$N^{ID}_{u_{i}}$ $>N^{text}_{u_{i}}$
where $N$ is the sequence length and they would be described in detail in section 3.4.
Once fed with $x^{ID}$ and $x^{text}$, RecLoRA begins its process. Our main predictor for labeling $y$ is the LLM
(8) $y$ $=LLM(x^{text})$
Thus, we need to fine-tune the pre-trained LLM with the recommendation dataset. To achieve memory efficiency, we choose to use LoRA for parameter-efficient tuning of the LLM, as introduced in Section 2.4. However, instead of using traditional LoRA, we design a Personalized LoRA Module to incorporate user-personalized knowledge. This module can be expressed as follows:
(9) $h_{j+1}$ $=Layer_{j}(h_{j})+LoRA(h_{j})$
$h_{j+1}$ $=Layer_{j}(h_{j})+PLoRA(h_{j},x^{ID})$
where $Layer_{j}$ is the $j_{th}$ layer of LLM, and PLoRA is our module, which takes the hidden states from the $j_{th}$ layer along with $x^{ID}$ as inputs and outputs the personalized representation. The details of PLoRA will be provided in Section 3.2.
3.2. Personalized Low-Rank Adaption
Intuitively, the most straightforward way for personalized low-rank adaptation is to maintain independent LoRA weights $(A_{u},B_{u})$ for each user $u$ separately, i.e.,
(10) $Y^{\prime}_{u}=XW+XA_{u}B_{u}^{T},\;\forall u\in\mathcal{U}.$
However, this approach is non-scalable, as its model complexity grows linearly with the number of users, leading to a tremendous amount of memory consumption. Moreover, it overlooks the overlap and commonality among behavior patterns of different users, which can lead to suboptimal performance.
Hence, to balance user-specific individual preference and inter-user pattern commonalities, we propose a Personalized LoRA Module. Specifically, as shown in Figure 3, we maintain a set of meta-LoRA weights $\{(A_{k},B_{k})\}_{k=1}^{N_{m}}$, which are designed to capture the diverse user behavior patterns. Each meta-LoRA weight $(A_{k},B_{k})$ can be regarded as a latent subspace to boost the expressiveness and capacity.
Then, we can obtain the personalized LoRA weights by combining them with a trainable gating network.
(11) $A_{p}B_{p}^{T}=\sum_{k=1}^{N_{m}}\alpha_{k}~{}A_{k}B_{k}^{T}.$
where $A_{p}B_{p}$ are the final personalized LoRA matrix, and $N_{m}$ is the number of meta-LoRA weights.
In this equation, the $k_{th}$ gating weight $\alpha_{k}$ represents the importance and relevance of the current user and the $k_{th}$ LoRA weight, acting as an indicator for personalization. Thus, To capture the complicated correlations and achieve personalized aggregation of the meta-LoRA weights, we introduce the conventional recommendation model (CRM) to provide personalized representations for the gating network. CRM takes discrete ID $x^{ID}$ as input and generates hidden states as personalized representations.
(12) $R_{c}=CRM(x^{ID})$
where $R_{c}\in\mathbb{R}^{n\times d_{c}}$ represents the CRM’s output. Specifically, we use a sequential behavior model, SIM(Pi et al., 2020b), as user history provides the most ideal personalization, and SIM is a state-of-the-art model for sequential behavior in recommendations.
After obtaining the representation, we adapt it to the gating weight to guide the meta-LoRA weights with personalized knowledge. We use a feed-forward network (FFN) with ReLU activation and LayerNorm as the adapter:
(13) $\beta$ $=W_{2}\times\text{ReLU}(\text{LN}(W_{1}R_{c}+b_{1})))+b_{2}$
and a softmax function to generate gating weights:
(14) $\alpha_{j}$ $=\frac{\exp(\frac{\beta_{j}}{tp})}{\sum_{j^{\prime}=1}^{N_{m}}\exp(\frac{\beta_{j^{\prime}}}{tp})}$
where $W_{1}\in\mathbb{R}^{d_{c}\times d_{h}}$, $W_{2}\in\mathbb{R}^{d_{h}\times d_{r}}$, $b_{1}\in\mathbb{R}^{d_{h}}$, $b_{2}\in\mathbb{R}^{d_{r}}$, and $\alpha\in\mathbb{R}^{n\times N_{m}}$ is the gating weight of Eq. (11). $tp$ is a temperature hyperparameter in the softmax function, making it smoother or sharper. Most importantly, softmax guarantees the distribution of personalized LoRA weights matches each meta-LoRA weights.
Finally, we obtain the personalized LoRA matrix by Eq. (11) and use it as the traditional LoRA method does in Eq. (9). In this way, this module dynamically generates one unique LoRA module personally for each user, realizing LoRA parameters personalization.
3.3. Few2Many Magnifier Strategy
During LLM fine-tuning, the ideal scenario would be exposing the entire recommendation dataset to the LLM, as studies have shown that the more data an LLM is exposed to, the better its performance. However, constraints on time and efficiency make it impractical to expose large datasets on the input side. Therefore, a more efficient approach is to inject large data knowledge without significantly increasing time costs. To achieve this, we propose the Few2Many Magnifier Strategy.
In the first stage, we train a conventional recommendation model using the full training dataset to obtain a well-trained model capable of providing sample-level personalized vector representations with high generalization based on the input ID modal data $x^{ID}$. This is achieved through a binary cross-entropy loss
(15) $\mathcal{L}^{ID}=\sum_{x_{i}^{ID}\in\mathcal{D}}[-y_{i}^{ID}\cdot\log(\sigma(\hat{y}_{i}^{ID}))-(1-y_{i}^{ID})\cdot\log(1-\sigma(\hat{y}_{i}^{ID}))].$
where $\sigma$ is sigmoid function, $y_{i}^{ID}$ is the true label and $\hat{y}_{i}^{ID}$ is the prediction label calculated in Eq. (3) Now CRM has fitted the full training space and learned most of the knowledge contained in the whole dataset.
In the second stage, we downsample the large-scale training set to obtain a smaller training set for the efficient parameter tuning of RecLoRA, which includes large language models. During this process, the fully trained traditional recommendation model from the first stage provides personalized information, ensuring full spatial perspectives. This information is combined to form a personalized LoRA matrix with sufficient generalization, as detailed in Section 3.2. Thus, even though the large language model sees only a small number of training samples during fine-tuning, it extends its receptive field to the full training space through the conventional recommendation models and personalized LoRA matrices. This greatly enhances the sample efficiency and recommendation performance of the large language model. It should be noted that during this process, only the personalized LoRA module is updated, while the traditional recommendation models and large language models remain fixed. The training loss used is the cross-entropy loss in traditional causal language modeling.
During the inference stage, large language model doesn’t directly give a pointwise score $\hat{y}^{text}\in\{0,1\}$. Therefore, we intercept the vocabulary scores, and then conduct a bidimensional softmax over the scores of binary key answer words. Specifically, the scores for “Yes” and “No” are $s_{y}$ and $s_{n}$ respectively. Then the pointwise scoring of LLMs can be written as:
(16) $\hat{y}^{text}=\frac{\text{exp}(s_{y})}{\text{exp}(s_{y})+\text{exp}(s_{n})}.$
This prediction will be used to calculate evaluation metrics.
3.4. Long-Short Modality Retriever
Intuitively, ideal personalized data involves a lifelong behavior sequence. However, the time cost escalates significantly with extended behavior sequences. To address efficiency issues while improving performance, we propose a Long-Short Modality Retriever.
It has been widely demonstrated that retrieval is effective in extracting important information from long behavior sequences for current sample prediction.(Qin et al., 2020; Pi et al., 2020b) However, the history length remains a trade-off between effectiveness and efficiency. Therefore, we retrieve long histories for input $x^{ID}$ in the CRM and short histories for input $x^{text}$ in the LLM. The CRM, trained on longer histories, can provide more sequential information and personalized knowledge for the LLM through LoRA parameters. This approach hardly increases the time cost, achieving a fantastic balance between effectiveness and efficiency, as will be shown in Section 4.5.
The retrieval method can be diverse and flexible. For example, we retrieve behaviors using semantic behavior encoding. Specifically, we input each behavior into the LLM and obtain the hidden states from the last layer of the LLM as its representation. We then apply principal component analysis (PCA) for dimension reduction and denoising. Finally, we calculate cosine similarities to identify the top-k most relevant behaviors to the current item.
4. Experiment
In this section, we conduct the experimental settings and results. Five research questions lead the following discussions, and our implementation code of RecLoRA is publicly available.
  • RQ1
    Does RecLoRA outperform existing baselines?
  • RQ2
    What are the influences of different components in RecLoRA?
  • RQ3
    What is the influence of meta-LoRA number $N_{m}$ in Eq. (11)?
  • RQ4
    How do the performance and time cost increase when extending a longer behavior sequence?
  • RQ5
    How does RecLoRA improve the sample efficiency for LLM?
Table 1. The dataset statistics.
Dataset #Users #Items #Samples #Fields #Features
MovieLens-25M 162,541 59,047 25,000,095 6 280,576
MovieLens-1M 6,040 3,706 970,009 10 16944
GoodReads 449,114 1,432,348 20,122,040 15 5,787,895
Table 2. The performance of different models in ID-based and LM-based settings. The best result is given in bold, and the second-best value is underlined. The symbol $\ast$ indicates a statistically significant improvement of RecLoRA over the best baseline with $p$-value ¡ 0.01.
Model MovieLens-25M MovieLens-1M GoodReads
AUC Log Loss Rel.Impr AUC Log Loss Rel.Impr AUC Log Loss Rel.Impr
ID-based DeepFM 0.8181 0.4863 3.52% 0.7978 0.5405 2.04% 0.7873 0.5027 1.47%
AutoInt 0.8138 0.4942 3.77% 0.7976 0.5398 2.06% 0.7866 0.5024 1.56%
DCNv2 0.8184 0.4905 3.50% 0.7977 0.5403 2.05% 0.7867 0.5022 1.55%
GRU4Rec 0.8169 0.4917 3.55% 0.7959 0.5423 2.29% 0.7869 0.5020 1.52%
Caser 0.8167 0.4917 3.39% 0.7952 0.5425 2.37% 0.7872 0.5023 1.49%
SASRec 0.8163 0.4892 3.54% 0.7984 0.5388 1.97% 0.7864 0.5022 1.59%
DIN 0.8258 0.4775 3.50% 0.7989 0.5389 1.90% 0.7880 0.4999 1.38%
SIM 0.8379 0.4664 1.59% 0.7992 0.5387 1.86% 0.7896 0.4993 1.18%
LM-based CTR-BERT 0.8079 0.5044 4.93% 0.7931 0.5457 2.64% 0.7176 0.5576 11.32%
TallRec 0.8324 0.4733 1.65% 0.7789 0.5580 4.51% 0.7759 0.5148 2.96%
ReLLa 0.8409 0.4662 1.22% 0.8005 0.5364 1.70% 0.7833 0.5065 1.99%
RecLoRA 0.8462 0.4552 - 0.8141 0.5248 - 0.7989 0.4916 -
Figure 4. The AUC performance of RecLoRA w.r.t. different meta-LoRA numbers $N_{m}$. RecLoRA manages to have already performed very well starting from $N_{m}=2$.
4.1. Experiment Setup
4.1.1. Datasets
We conduct experiments on three real-world datasets (i.e., GoodReads [^1] https://mengtingwan.github.io/data/goodreads.html, MovieLens-1M [^2] https://grouplens.org/datasets/movielens/1m/ and MovieLens-25M [^3] https://grouplens.org/datasets/movielens/25m/). We show the dataset statistics in Table 1 and give detailed data preprocessing information in Appendix A due to page limitations.
Table 3. The performance of different variants of RecLoRA. We remove different components of RecLoRA to evaluate the contribution of each part to the model. The best result is given in bold, and the second-best value is underlined..
Model ML-25M ML-1M GoodReads
AUC Log Loss Rel.Decr AUC Log Loss Rel.Decr AUC Log Loss Rel.Decr
RecLoRA 0.8462 0.4552 - 0.8141 0.5248 - 0.7989 0.4916 -
RecLoRA(w/o meta-LoRA) 0.8360 0.4662 -1.21% 0.7869 0.5517 -3.34% 0.7740 0.5181 -3.12%
RecLoRA(w/o meta-LoRA same param) 0.8370 0.4759 -1.09% 0.7820 0.5548 -3.94% 0.7735 0.5125 -3.17%
RecLoRA(w/o long retriver) 0.8501 0.4761 +0.46% 0.8026 0.5321 -1.41% 0.6744 0.6231 -15.6%
RecLoRA(w/o short retriver) 0.8473 0.4606 +0.13% 0.8001 0.5359 -1.71% 0.7971 0.4924 -0.22%
RecLoRA(w/o long-short retriver) 0.8453 0.4787 -0.11% 0.7978 0.5373 -2.00% 0.6882 0.6073 -13.9%
Table 4. The performance and time cost of ReLLa and RecLoRA in GoodReads w.r.t. different hist length in text and id inputs.
Hist Length AUC Training Inference
Text ID Value Improve Time Increase Time Increase
10 - 0.7776 - 1.4s/sample - 0.6s/sample -
30 - 0.7828 +0.6% 3.34s/sample +138.78% 1.06s/sample +76.67%
60 - 0.7833 +0.73% 6.41s/sample +358.16% 1.83s/sample +205.55%
10 30 0.7822 +0.59% 1.42s/sample +2.04% 0.63s/sample +5.55%
10 60 0.7989 +2.73% 1.45s/sample +4.08% 0.65s/sample +8.33%
4.1.2. Evaluation Metrics
To assess the effectiveness of our methods, we use AUC (Area Under the ROC Curve), Log Loss (binary cross-entropy loss), and Rel.Impr (Relative Improvement) as evaluation metrics. In the context of CTR prediction, even minor variations, such as a 0.001 increase in AUC or a 0.001 decrease in Log Loss, are considered significant improvements (Lian et al., 2018; Wang et al., 2021b).
4.1.3. Baseline Models
We have divided our baselines into two main categories:(1) traditional ID-based, containing DeepFM (Guo et al., 2017b), AutoInt (Song et al., 2019), and DCNv2 (Wang et al., 2021b) as exemplars of feature interaction models and GRU4Rec (Hidasi et al., 2016), Caser (Tang and Wang, 2018), SASRec (Kang and McAuley, 2018b), DIN (Zhou et al., 2018), and SIM (Pi et al., 2020b) as key representatives of user behavior models (2) LM-based, containing CTR-BERT(Muhamed et al., 2021a), TALLRec (Bao et al., 2023), and ReLLa (Lin et al., 2024b). Detailed description of these baselines are provided in Appendix B.
4.1.4. Implementation Details
We have provided our details such as base-model, hyperparameters and prompts in Appendix C.
4.2. Overall Performance (RQ1)
We evaluate the performance of RecLoRA in comparison to existing baseline models, with the results reported in Table 2. It is important to note that most of the other recommendation baseline models are trained in full-shot settings with the entire training set, while only TallRec, ReLLa and RecLoRA are trained on few-shot training sets with 70000(¡10%) training samples in all three datasets. We set the length of text and ID behavior sequences to 10 and 60 respectively in all three datasets. ReI.Impr denotes the relative AUC improvement rate of RecLoRA against each baseline. Our observations from Table 2 are as follows.
  • SIM achieves the best performance among all the ID-based baseline models. By applying user behavior retrieval, SIM effectively reduces the noise in user sequences, which is fundamentally beneficial for CTR prediction. Additionally, the traditional LM-based CTR model CTR-BERT perform worse than most ID-based traditional CTR models, consistent with the results reported in [40, 65].It only incorporate the small language model BERT [2] for pure text-based recommendations, resulting in inferior performance.
  • ReLLa generally achieves performance comparable to the best ID-based baseline model, SIM. ReLLa incorporates retrieval-enhanced instruction tuning (ReiT) and utilizes only 10% of the training data. This observation demonstrates the effectiveness of both tuning LLM and retrieval in the input sequences.
  • RecLoRA outperforms the state-of-the-art baselines significantly with a p-value ¡ 0.01 against the best baseline, validating the effectiveness of our proposed Personalized LoRA module. The better performance with only 10% of the training data for tuning LLM demonstrates that our RecLoRA model effectively captures user interest and achieves superior user personalization.
4.3. Ablation Study (RQ2)
In this section, we conduct ablation experiments to analyze the effectiveness of the main components in RecLoRA: the Personalized LoRA module and Long-Short Modality Retriever.
  • RecLoRA (Ours): This is the complete version of our proposed method.
  • RecLoRA (w/o meta-LoRA): We remove the Personalized LoRA module and only maintain a LoRA module with a rank of 8.
  • RecLoRA (w/o meta-LoRA same param): Since the personalized meta-LoRA module has $N_{m}=16$ times more training parameters, we maintain the parameter count by replacing the meta-LoRA weights with a single LoRA weight, increasing the LoRA rank from 8 to 128.
  • RecLoRA (w/o long retriever): To evaluate the effect of retrieval history in $x^{ID}$, we remove the long retriever module and only use the recent history.
  • RecLoRA (w/o short retriever): To evaluate the effect of retrieval history in $x^{text}$, we remove the short retriever module and only use the recent history.
  • RecLoRA (w/o long-short retriever): To evaluate the overall effect of retrieval history, we remove both the long and short retrievers, using only the recent history.
The results are shown in Table 3. We observe that the performance of RecLoRA significantly decreases when the personalized LoRA module is removed. This finding confirms that the lack of user personalization knowledge in the LLM tuning process leads to suboptimal performance. Our proposed personalized LoRA module effectively addresses this issue and improves performance. Additionally, we observe from the third line that the improvement is not merely due to an increase in the parameter count by $N_{m}$ times.
Meanwhile, we observe that removing either the long or short retriever generally results in a performance drop. This highlights the importance of considering lifelong user sequences and demonstrates that input from different modalities both benefit from the retrieval of long sequences.
Figure 5. The performance of RecLoRA w.r.t. different hist length in id inputs.
4.4. Hyperparameter Study (RQ3)
In this section, we conduct hyperparameter experiments to analyze the influence of the meta-LoRA number $N_{m}$ in personalized LoRA module. The results are shown in Figure 4.
As shown in Figure 4, the evaluation metric AUC generally increases with the hyperparameter $N_{m}$, reaching a maximum value at $N_{m}=16$. This experiment demonstrates that increasing the parameter count improves overall performance, as the hyperparameter $N_{m}$ determines the number of meta-LoRA weights and the associated trainable parameters.
Furthermore, when $N_{m}=1$ , the personalized LoRA module collapses into a single LoRA module without CRM instruction, resulting in the worst performance. Notably, even starting from $N_{m}=2$, RecLoRA performs very well, even comparable to $N_{m}=16$ in GoodReads. This illustrates the effectiveness of personalized LoRA without significantly increasing the parameter count.
4.5. Time Efficiency (RQ4)
In this section, we conduct time efficiency experiments to analyze the balance of performance and time cost when increasing sequence lengths in the RecLoRA model. The results of GoodReads are shown in Figure 5 and Table 4. And The results of the other two datasets are similar and omitted in Table 4 due to page limits.
Figure 5 displays the performance of RecLoRA as ID sequence lengths increase. We observe that the performance improves significantly with longer ID sequences. Since SIM provides instructions for meta-LoRA weights aggregation in the personalized LoRA module, better performance of SIM theoretically enhances RecLoRA’s performance, which is consistent with our experimental results.
Table 4 shows the performance and time cost of RecLoRA with different historical lengths in text and ID inputs. It reveals that while increasing text sequence lengths improves performance by +0.73%, it also leads to a substantial and unacceptable increase in training and inference time costs of +358.16% and +205.55%, respectively. In contrast, not only the performance is improved significantly by +2.73% when extending ID history lengths, but also the corresponding training and inference time only increase by +4.04% and +8.33% respectively, which are negligible compared to the time increases associated with longer text sequences. These results highlight RecLoRA’s advantage in balancing effectiveness and efficiency.
Figure 6. The AUC performance of ReLLa and RecLoRA models w.r.t different sample numbers used to tune LLM. RecLoRA has obviously a higher sample efficiency compared to ReLLa.
4.6. Sample Efficiency (RQ5)
In this few-shot setting, we want to explore how much data is necessary for LLMs to adapt to CTR distributions. So we investigate the sample efficiency by varying the number of samples used in training LLM, following previous works(Lin et al., 2024b). In Figure 6, we report the AUC performance of ReLLa (a few-shot baseline) and RecLoRA with different sample sizes (ranging from 2,048 to 70,000). The text and ID behavior sequence lengths are set to 10 and 60, respectively.
As depicted in Figure 6, both ReLLa and RecLoRA show performance enhancement as the number of samples gradually increases. However, with the same sample number, RecLoRA consistently outperforms ReLLa by a significant margin. Moreover, even with a small sample size (10,000), RecLoRA performs better than ReLLa with a relatively large sample size (70,000). This is due to RecLoRA’s Few2Many Strategy, which utilizes a fully pre-trained CRM for instruction, helping LLMs adapt to CTR distributions better with fewer samples.
With a limited number of training samples, RecLoRA demonstrates remarkable sample efficiency and considerable few-shot inference ability, benefiting from the mutual assistance of the open-world knowledge of LLMs and the collaborative signals of CRMs.
5. Related Works
5.1. Large Language Model for Recommendation
Previous work (Lin et al., 2023a) has suggested that the employment of language models to recommender systems can generally be categorized based on the roles they play in the recommendation pipeline. i.e., feature engineering (Liu et al., 2023a; Borisov et al., 2022; Li et al., 2023d; Mysore et al., 2023; Carranza et al., 2023; Christakopoulou et al., 2023), feature encoder (Muhamed et al., 2021b; Hou et al., 2022; Yu et al., 2021; Wang et al., 2022; Hou et al., 2023a; Zhang et al., 2022; Fu et al., 2023; Yuan et al., 2023; Qiu et al., 2021; Li et al., 2023c; He et al., 2022; Muhamed et al., 2021b), scoring/ranking function (Liu et al., 2022; Kang et al., 2023; Zhang et al., 2021; Li et al., 2023f; Bao et al., 2023; Li et al., 2023e; Zhang and Wang, 2023; Mao et al., 2023; Hua et al., 2023a; Geng et al., 2023; Hua et al., 2023b; Zhang et al., 2023a; Hou et al., 2023b; Chen, 2023; Petrov and Macdonald, 2023; Wang and Lim, 2023).
In feature engineering, large language models (LLMs) process raw data (e.g., user profiles and item descriptions) as input, and output open-world knowledge or potential new attributes for data augmentation with carefully designed prompts or templates. For example, KAR (Xi et al., 2023) utilizes the potential knowledge of user preferences and item attributes by requesting LLMs with factorization prompting techniques. GENRE (Liu et al., 2023a) employs LLMs to generate news summarization, synthetic pieces, and user profiles. The obtained knowledge serves as augmented features and improves performances of recommenders in a model-agnostic manner.
In feature encoders, LLMs are used as auxiliary textual encoders to both enrich the user/item representations with semantic information and enable cross-domain recommendation with the open-world natural language interface. For instance, U-BERT (Qiu et al., 2021) enhances user representation by encoding review texts into dense vectors via BERT. UniSRec (Hou et al., 2022) and VQ-Rec (Hou et al., 2023a) apply a fixed BERT as the encoder for item descriptive texts, to achieve unified cross-domain sequential recommendation.
In scoring/ranking function, instead of doing assistant tasks for recommendation (e.g., feature engineering or feature encoder), LLMs are adopted to do the scoring or ranking task, which is the core component of recommendation. In this case, LLMs try to accomplish either the item scoring task (Liu et al., 2022; Kang et al., 2023; Zhang et al., 2021; Li et al., 2023f; Bao et al., 2023; Li et al., 2023e; Zhang and Wang, 2023; Mao et al., 2023; Lin et al., 2024b), or item generation task (Hua et al., 2023a; Geng et al., 2023; Hua et al., 2023b; Zhang et al., 2023a; Hou et al., 2023b; Chen, 2023; Petrov and Macdonald, 2023; Wang and Lim, 2023). Also, various works (Geng et al., 2022; Cui et al., 2022; Zhang et al., 2023c; Liu et al., 2023b; Sun et al., 2023; Dai et al., 2023) attempt to explore the multi-task capacity of LLMs, and instruct LLMs with various ways to solve the multiple tasks (e.g., both scoring and generation) through a unified language interface.
5.2. Long Sequence User Modeling
During recommendation task, Kim (Kim and Chan, 2003) argues that long-term sequence mean general interest, which is back to one’s mind and important for personalization. Existing approaches for addressing long-term user sequence mainly focus on memory network and retrieval methods. Hierarchical Periodic Memory Network(HPMN) (Ren et al., 2019) proposes a hierarchical and periodical updating mechanism for capturing multi-scale sequential user interests. The Memory Augmented Neural Network (MIMN) (Pi et al., 2019) stores behaviors in a memory matrix at the user interest center and updates the memory for new users. Sequential Interest Modeling (SIM) (Pi et al., 2020a) and User Behavior Retrieval for CTR (UBR4CTR) (Qin et al., 2020) have introduced retrieval-enhanced history and two-stage frameworks to catch user patterns in the past which are related to current targets. UBR4CTR employs BM25 as the relevance metric in its initial stage. In contrast, the original SIM has two variations with different designs. SIM Hard selects relevant items based on the same category as the target item, while SIM Soft utilizes the inner product of pre-trained item embeddings to determine relevance. Efficient Temporal Attention (ETA) (Chen et al., 2021) employs locality-sensitive hashing (LSH) to encode item embeddings and retrieve relevant items from long sequences via Hamming distance. Sequential Dynamic Interest Modeling (SDIM) (Cao et al., 2022) samples behavior items that share the same hash signature as the target item through multiple rounds of hash collisions, and then linearly aggregates these sampled behavior items to capture user interests.
6. Conclusion
In this paper, we explore the application of large language models in recommendation systems. We begin by addressing the current challenges related to personalization and long sequences. To tackle these issues, we propose RecLoRA, a model designed for lifelong personalized low-rank adaptation. RecLoRA includes a Personalized LoRA module, which maintains distinct LoRAs for individual users. Additionally, a Long-Short Modality Retriever extracts varying history lengths for different modalities, enhancing performance with minimal time cost. We also employ a conventional recommendation model to extend the scope of LLMs to the full training space. RecLoRA shows promising performance in offline experiments on three public datasets. Further, ablation, hyperparameter, and efficiency studies demonstrate its strong balance between effectiveness and efficiency. For future work, we plan to explore more advanced structures for personalization in LLMs and address the fairness issue, which is crucial for achieving complete personalization.
Appendix A Data Preprocessing
Our experiments are conducted on three real-world public datasets (i.e., MovieLens-25M, MovieLens-1M, GoodReads), and the statistics of the processed datasets are show in Table 1. All these datasets are split into training, validing and testing sets with ratio of 8:1:1 according to the global timestamp (Qin et al., 2021).
  • MovieLens-25M has a scoring range from 0 to 5, with increments of 0.5. We label samples with ratings above 3.0 as positive, and the rest as negative.
  • MovieLens-1M contains user-movie integer ratings ranging from 0 to 5. Samples with ratings of 4 and 5 are labeled as positive and the rest as negative.
  • GoodReads is a book recommendation dataset from GoodReads website with ratings ranging from 1 to 5. We transform the ratings into binary labels with a threshold of 4.
We apply 5-core filtering to ensure each user or item has at least five interaction records.
Under the few-shot setting with a particular number of training data, we uniformly sample $N$ data instances from the training set, which is then fixed during few-shot tuning.
Appendix B Baseline Implementation
In this section, we describe the baseline models. Baseline models can broadly be divided into two main categories: (1) traditional CTR models, which primarily use one-hot encoded IDs as inputs, and (2) LM-based models, which integrate pre-trained language models to approach CTR prediction as either a text classification or a sequence-to-sequence problem.
Within traditional CTR models, we distinguish between (1) feature interaction models and (2) user behavior models. We have selected DeepFM (Guo et al., 2017b), AutoInt (Song et al., 2019), and DCNv2 (Wang et al., 2021b) as exemplars of feature interaction models, and GRU4Rec (Hidasi et al., 2016), Caser (Tang and Wang, 2018), SASRec (Kang and McAuley, 2018b), DIN (Zhou et al., 2018), and SIM (Pi et al., 2020b) as key representatives of user behavior models. For the feature interaction models, we implement average pooling across users’ historical behaviors and treat the results as additional feature fields.
SIM is a classic sequential CTR model that employs user behavior retrieval techniques to enhance recommendation performance. We include it in our evaluation to ensure a comprehensive and fair comparison, maintaining consistency with the retrieval method used in RecLoRA.
For LM-based CTR models, we have chosen CTR-BERT (Muhamed et al., 2021a), TALLRec (Bao et al., 2023), and ReLLa (Lin et al., 2024b) as baseline models to represent this category, and they are representatives of traditional language models and large language models respectively.
B.1. Traditional CTR Models
We choose the embedding size from {32, 64} on three datasets. The dropout rate is selected from {0.0, 0.1, 0.2}. The activation function is fixed to ReLU. The learning rate is selected from $1\times 10^{-3},5\times 10^{-4},1\times 10^{-4}$ and AdamW (Loshchilov and Hutter, 2017) optimizer is used. The batch size is selected from {64,128,256, 512,1024}. More model-specific hyperparameter settings are shown as follows:
  • DeepFM (Guo et al., 2017b). We choose the size of DNN layer from {128, 256} and the number of DNN layers from {3, 6, 9, 12}.
  • AutoInt (Song et al., 2019). The attention layers is selected from {3, 6, 9, 12} and the attention size is selected from {64, 128, 256}. The number of attention heads are all set to 1.
  • DCNv2 (Wang et al., 2021b). We choose the size of DNN layers from {128, 256} and the number of DNN layers and cross layers are from {3, 6, 9, 12}.
  • GRU4Rec (Hidasi et al., 2016). The number of GRU layers is selected from {1, 2, 3}. The GRU hidden size and DNN hidden size is selected from {64, 128, 256}.
  • Caser (Tang and Wang, 2018). The number of vertical convolution kernels is selected from {2, 4, 8}. The number of horizontal convolution kernels is selected from {4, 8, 16}. The number of DNN layers is selected from {1,2,3}. The DNN hidden size is selected from {64, 128, 256}.
  • SASRec (Kang and McAuley, 2018b). The number of attention heads is selected from {1, 2, 4}. The number of attention layers is selected from {1, 2, 3}. The attention size is selected from {64, 128, 256}. The number of DNN layers is selected from {1,2,3}. The DNN hidden size is selected from {64, 128, 256}.
  • DIN (Zhou et al., 2018). The number of DIN attention layers and DNN layers are selected from {1, 2, 3}. The DNN hidden size is selected from {64, 128, 256}.
  • SIM (Pi et al., 2020b). The number of attention layers and DNN layers are selected from {1, 2, 3}. The DNN hidden size is selected from{64, 128, 256}.
B.2. LM-based Models
The structure of the pre-trained language models is kept unchanged. And AdamW (Loshchilov and Hutter, 2017) optimizer is used for all the baselines. The detailed training settings are as follows:
  • CTR-BERT (Muhamed et al., 2021a). We maintain a two-tower model structure based on the BERT (Devlin et al., 2018) model to encode the user and item information respectively. The total number of tuning epochs is set to 10. The batch size is set to 1024. The learning rate is set to $5\times 10^{-5}$ with linear decay. The warmup ratio is 0.05.
  • TallRec (Bao et al., 2023) is a supervised fine-tuning framework with Llama as a backbone pretrained language model for recommendation task. For fair comparison, we generate the same history in text as ID-based baselines. The total number of tuning epochs is selected from {10,5}. The batch size is selected from {64,128,256}. The learning rate is selected from $1\times 10^{-3},5\times 10^{-4},1\times 10^{-4}$}.
  • ReLLa (Lin et al., 2024b) uses user behavior retrieval and supervised fine-tuning with Llama as a backbone pretrained language model for recommendation task. For fair comparison, the retrieval methods in SIM, ReLLa and RecLoRA maintain the same. The total number of tuning epochs is selected from {10,5}. The batch size is selected from {64,128,256}. The learning rate is selected from $1\times 10^{-3},5\times 10^{-4},1\times 10^{-4}$}
Appendix C Implementation Detals
We selected Vicuna-7B (Chiang et al., 2023), provided by FastChat [^4] https://github.com/lm-sys/FastChat, as the foundational large language model (LLM) for our experiments. All computations were performed using V100 GPUs. To enhance training resource efficiency, we employed 8-bit quantization. The LoRA configuration was set with a rank of 8, an alpha value of 16, and dropout set to 0.0. LoRA update matrices were specifically applied to the query and value projection matrices within the attention blocks.
For the instruction tuning process, our detailed configuration such as optimizer and hyperparameters are provided in we utilized the AdamW (Loshchilov and Hutter, 2017) optimizer, configuring it with zero weight decay. The training process was carried out with batch sizes chosen from $\{64,128,256\}$ and initial learning rates selected from $\{1\times 10^{-3},5\times 10^{-4},3\times 10^{-4},1\times 10^{-4}\}$ using a linear scheduler. For dataset-specific training, the maximum number of epochs was set to 10 for the MovieLens-25M datasets and 5 for MovieLens-1M and GoodReads datasets. Detailed configurations of the baseline models are provided in Appendix B. Additionally, the hard prompt templates for textual input-output pairs and item descriptions across all datasets are detailed in Figure 2
Furthermore, when formulating the hard prompt template for RecLoRA, we excluded all pure ID fields—namely, User ID, Book ID fields in GoodReads, User ID, Movie ID, and Zipcode fields in the MovieLens-1M dataset, as well as User ID and Movie ID fields in the MovieLens-25M dataset. This exclusion was based on findings that LLMs exhibit limited perceptual capabilities for pure ID texts (Lin et al., 2023a). In contrast, other fields were utilized as user profile or item information in the prompts, as outlined in Section 2.3 and Figure 2. Notably, for other CTR baseline models, we retained all feature fields and user behavior sequences as inputs, ensuring no features were discarded.